# Spontaneous emission of a three-level artificial atom in a one-dimensional open waveguide

### v1

@@ IV.1 Solution in real-space
Let’s assume that initially there are no photons in a waveguide, and a three-level atom is in upper excited state $\left|f\right\rangle$:
@@
In this case we have the process of spontaneous decay, which due to three levels results in emission of two photons.
Using initial condition (17) in (16), (15) and (10), we find exact expressions for all three probability amplitudes of a wavefunction:
@@
where composite theta-function:
@@
Since here we work in a chiral approximation, all fields moves only right, so the emission of spontaneous photons only happens in forward direction, $x_{1,2}\equiv x>0$.
@@ IV.2 State detection probabilities
Probability amplitudes (18), beside $\alpha(t)$, contain photonic dependence, proportional to real-space coordinate $x$ where photon is presumably located. So we can treat square root of those amplitudes as density of probability to find a photon in certain waveguide region. To find full probabilities to locate desired system state, one needs to integrate it over entire waveguide. By doing so, we obtain:
@@
Figure 2: Probabilities to detect a certain state in entire waveguide, plotted with expressions (20): red solid line - transmon fully excited in level $\left|f\right\rangle$ and no photons presented; blue dashed line - transmon in first-excited state $\left|e\right\rangle$ and single photon in a waveguide; green dash-dotted line - transmon in a ground state $\left|g\right\rangle$, and two photons in a waveguide. Sum of all probabilities is equals to one and shown by thin dotted line. Different ratios of $\Gamma_{2}/\Gamma_{1}$ are shown above plots. Time is measured in $\Gamma_{2}$ units. Decay of upper level is constant, $\Gamma_{2}/\Omega_{1}=0.02$.
As one can see from (20), probabilities to detect a certain state of wavefunction (3) are independent on atomic resonant frequencies, and contain only decay rates $\Gamma_{1}$ and $\Gamma_{2}$. Plots of state-detection probabilities for a few $\Gamma_{2}/\Gamma_{1}$ ratios are presented in Fig. 2.
The decay of upper level $\left|f\right\rangle$ is not dependent on $\Gamma_{1}$, i.e. it’s stays the same for different $\Gamma_{2}/\Gamma_{1}$ ratios since we choose to take $\Gamma_{2}/\Omega_{1}=0.02$ constant (red solid line if Fig. 2). However, it is not true for the decay with an emission of single or two photons in the entire waveguide. There are two extreme cases: the decay of upper level is much larger, that of the middle one, $\Gamma_{2}/\Gamma_{1}\gg 1$, and inverse picture, when $\Gamma_{2}/\Gamma_{1}\ll 1$. For first case, the decay of upper levels is so fast, that transmon after short period of time starts to behave somewhat similar to a regular two-level system, formed by states $\left|e\right\rangle$ and $\left|g\right\rangle$ (see blue and green lines in Fig. 2a). For $\Gamma_{2}/\Gamma_{1}\leftarrow\infty$ first-excited state $\left|e\right\rangle$ become metastable. In second case decay of upper level $\left|f\right\rangle$ is almost instantly followed by decay of middle level $\left|e\right\rangle$, so, despite the cascaded process of emission, the two photons are produced nearly simultaneously (blue line in Fig. 2d is very close to zero, and green line is close to unity).
Typical transmon circuits have coupling ratio between first-excited and second-excited states as $g_{2}(k)/g_{1}(k)\sim\sqrt{3/2}$, see expressions (3.3-3.4) in Koch07 . The corresponding ratio of decay rates thus is $\Gamma_{2}/\Gamma_{1}\sim 1.5$, although some works mention very similar value of $\Gamma_{2}/\Gamma_{1}\sim 2$ Gasparinetti17 . We plot probabilities for common relation of $\Gamma_{2}/\Gamma_{1}=1.5$ in Fig. 2b. Albeit, recent progress in superconducting quantum circuits made possible creation of the so-called giant-atom transmon with highly tunable ratio $\beta=\Gamma_{2}/\Gamma_{1}\approx 0.3-60$ Vadiraj21 . However, further down we stick to more common transmon types with ratios of $\Gamma_{2}/\Gamma_{1}=1.5$ and $\Gamma_{2}/\Gamma_{1}=3$.
@@ IV.3 Two-photon spectrum
By using Fourier transform (7), we can translate the real-space photon amplitudes (10) and (15) into a k-space of frequency modes:
@@
where in $\beta(k,t)$ photon frequency $\omega$ corresponds to first emitted photon.
Note that if we take $t\rightarrow\infty$, single photon amplitude (21), as well as $\alpha(t)$, will be equals to zero. However, two-photon amplitude (22) for large times takes non-zero value. It is logical prediction, because final state of excited three-level system is obviously a two photons in a waveguide with the transmon being in the ground state.
Let’s define the spectral density of spontaneous emission of initially excited transmon as follows:
@@
We plot this two-photon density in Fig. 3.
Unlike the probabilities (20), spectral density (23) depends not only on decay rates $\Gamma_{1},\Gamma_{2}$, but also on resonant frequencies of 3LE. It is typical to describe energy-level spacing of a transmon via an anharmonicity parameter (Koch07, ):
@@
where $E_{J}$ is a Josephson energy, and $E_{C}$ is a charging energy - common characteristics of superconducting quantum circuits Blais21 . Subscript $r$ stands for ’relative’, thus expression (24) is dimensionless and shows relation of energies between states $\left|f\right\rangle,\;\left|e\right\rangle$ and energies between levels $\left|e\right\rangle,\;\left|g\right\rangle$. In literature it is also common to use a non-relative anharmonicity $\alpha=(\Delta_{\Omega}-\Omega_{1})=\alpha_{r}\Omega_{1}$, that is measured in hertz, but here we choose to use only relative parameter to avoid confusion with probability amplitude $\alpha(t)$. Typical experimental values of $\alpha$ usually falls within the range 100-300 MHz Krantz19 , or in relative units from $2\%$ up to $6\%$. Note that anharmonicity for superconducting qubits are always negative, due to smaller spacing between upper levels.
@@
Figure 3: Two-photon spectral density of spontaneous emission from three-level transmon $S_{sp}(\omega_{1},\omega_{2})\Gamma_{2}^{2}$, plotted using expression (23). (a) high anharmonicity $\alpha_{r}=-5\%$ and typical decay rates relation $\Gamma_{2}/\Gamma_{1}=3/2$; (b) smaller anharmonicity $\alpha_{r}=-3\%$, and $\Gamma_{2}/\Gamma_{1}=3/2$; (c) faster decay of upper level $\Gamma_{2}/\Gamma_{1}=3$ and $\alpha_{r}=-3\%$; (d) same decay $\Gamma_{2}/\Gamma_{1}=3$ and smaller anharmonicity $\alpha_{r}=-2\%$. Decay of upper level taken constant $\Gamma_{2}/\Omega_{1}=0.02$, except for (e), where weak-coupling is analyzed for $\Gamma_{2}/\Omega_{1}=0.001$, $\Gamma_{2}/\Gamma_{1}=1.5$ and $\alpha_{r}=-3\%$.
Figure 4: Spectral density of two identical photons (25) vs frequency detuning - (a) and (b), and vs detuning and anharmonicity - (c) and (d). (a) and (c) are plotted for $\Gamma_{2}/\Gamma_{1}=3/2$ ratio, and (b) and (d) for $\Gamma_{2}/\Gamma_{1}=3$. In (a), (b) solid magenta line corresponds to $\alpha_{r}=-3\%$, green dashed line to $\alpha_{r}=-4\%$, and black dotted line to $\alpha_{r}=-6\%$. Vertical grey lines show points for $\delta_{\omega}/\Omega_{1}=\alpha_{r}/2$. In (c), (d) thin white lines indicate resonant frequency $\omega=\Omega_{1}$. Decay of upper level taken constant $\Gamma_{2}/\Omega_{1}=0.02$.
@@
From Fig. 3 we see that spectral density of spontaneous emission have two distinct peaks, which are a result of non-distinguishability of photons. Each peak corresponds to one resonant photon with frequency $\omega=\Omega_{1}$ and second photon with smaller frequency $\omega=(1+\alpha_{r})\Omega_{1}$ (remember that $\alpha_{r}$ is negative). But because both photons are indistinguishable, we can’t exactly say which photon is ’first’, and which is ’second’, thus two peaks appear. For bigger anharmonicity, two peaks are further apart, because of change of transition frequencies in 3LE. If we implement weak-coupling, $\Gamma/\Omega_{1}=0.001$ (in practical circuits its usually $\Gamma<5$ MHz), both peaks are tighten into almost point-like shapes (see Fig. 3e). That means that we have two photons with distinct frequencies, correspondent to atom-transitions, $\Omega_{1}$ and $(1+\alpha_{r})\Omega_{1}$. This behavior is in good agreement with experimental work by Gasparinetti et al Gasparinetti17 .
@@
However, if we take stronger coupling, typical for circuit QED systems (for example, $\Gamma_{2}\sim 100$ MHz), we see the appearance of frequency correlation between the two photons. This is expressed by the appearance of third peak at resonant frequency $\omega/\Omega_{1}=1$, which is start to become visible for $\Gamma_{2}/\Gamma_{1}=1.5$ in Fig. 3(b), and more pronounced for $\Gamma_{2}/\Gamma_{1}=3$ in Fig. 3(c)-(d). Thus, even if three-level system have non-symmetrical level structure, there is a probability that during spontaneous decay emitted photons would have same frequency $\omega_{1}=\omega_{2}=\Omega_{1}$. This effect is connected to interference between first emitted photon and 3LE being still excited in state $\left|e\right\rangle$, which is arise due to almost equidistant level structure of low-anharmonicity transmon and rather strong coupling between the artificial atom and the field modes.
@@ IV.4 Emission spectra for identical photons
Let’s study spectral density (23) for identical photons in closer detail. Taking $\omega_{1}=\omega_{2}$ we get:
@@
where we introduce detuning from resonant frequency for both photons as $\delta_{\omega}=\omega-\Omega_{1}$.
Function (25) have two maximums: first one corresponds to frequency between lower levels $\Omega_{1}$ and thus $\delta_{\omega}=0$, and second one corresponds to decay of upper level with anharmonicity shift, $\delta_{\omega}=-\alpha_{r}\Omega_{1}/2$. Both these peaks can be observed in Fig. 4(a)-(b), for decay rates $\Gamma_{2}/\Gamma_{1}=3/2$ and $\Gamma_{2}/\Gamma_{1}=3$. Two-dimensional plot of (25), showing how spectral density also changes with anharmonicity, presented in Fig. 4(c)-(d).
The maximum of spectral density for identical photons should be for $\alpha_{r}\rightarrow 0$, i.e. in the case of equidistant levels with $\Omega_{2}=2\Omega_{1}$. However, as can be seen from Fig. 4, frequencies of two photons not always have to be resonant with $\Omega_{1}$. For example, if ratio between decay rates is $\Gamma_{2}/\Gamma_{1}=1.5$, like in Fig. 4(a), emitted photons both would have maximum spectral density on frequency, close to $\Omega_{1}(1+\alpha_{r}/2)$. This also can be observed on two-frequency graph Fig. 3(b) in diagonal cross-section. Thus, by changing both decay rates $\Gamma_{1},\Gamma_{2}$ and anharmonicity parameter $\alpha_{r}$ of excited three-level transmon, it is possible to create correlated photon pairs with same frequencies. Superconducting circuits are very well suitable for this task, since they allow tuning of couplings and anharmonicity parameters either by sample production techniques or via application of external magnetic fields.

### v2

@@ IV.1 Solution in real-space
Let’s assume that initially there are no photons in the waveguide, and a three-level atom is in the upper excited state $\left|f\right\rangle$:
@@
In this case we have the process of spontaneous decay, which, due to the three levels, results in the emission of two photons.
Using the initial condition (17) in (16), (15) and (10), we find the exact expressions for all three probability amplitudes of the wavefunction:
@@
where the composite theta-function is:
@@
Working within the chiral approximation implies that all fields are right-moving. Consequently, spontaneous photon emission is restricted to the forward direction along the positive x-axis, $x_{1,2}\equiv x>0$, since we place transmon in the center of the waveguide $x=0$.
@@ IV.2 State detection probabilities
Figure 2: Probabilities to detect a certain state in the entire waveguide, plotted with expressions (20): red solid line — transmon fully excited in level $\left|f\right\rangle$ and no photons present; blue dashed line — transmon in the first-excited state $\left|e\right\rangle$ and a single photon in the waveguide; green dash-dotted line — transmon in the ground state $\left|g\right\rangle$, and two photons in the waveguide. The sum of all probabilities is equal to one and shown by a thin dotted line. Different ratios of $\Gamma_{2}/\Gamma_{1}$ are shown above the plots. Time is measured in $\Gamma_{2}$ units. The decay rate of the upper level is constant, $\Gamma_{2}/\Omega_{1}=0.02$.
Probability amplitudes (18), besides $\alpha(t)$, contain a photonic dependence proportional to the real-space coordinate $x$ where the photon is presumably located. So we can treat the square root of those amplitudes as the density of probability to find a photon in a certain waveguide region. To find the full probabilities to locate a desired system state, one needs to integrate it over the entire waveguide. By doing so, we obtain:
@@
As one can see from (20), the probabilities to detect a certain state of the wavefunction (3) are independent of the atomic resonant frequencies and contain only the decay rates $\Gamma_{1}$ and $\Gamma_{2}$. Plots of state-detection probabilities for a few $\Gamma_{2}/\Gamma_{1}$ ratios are presented in figure 2.
The decay of the upper level $\left|f\right\rangle$ does not depend on $\Gamma_{1}$, i.e., it stays the same for different $\Gamma_{2}/\Gamma_{1}$ ratios since we choose to keep $\Gamma_{2}/\Omega_{1}=0.02$ constant (red solid line in figure 2). However, this is not true for the decay with the emission of a single or two photons in the entire waveguide. There are two extreme cases: the decay of the upper level is much faster than that of the middle one, $\Gamma_{2}/\Gamma_{1}\gg 1$, and the inverse picture, when $\Gamma_{2}/\Gamma_{1}\ll 1$. For the first case, the decay of the upper level is so fast that the transmon, after a short period of time, starts to behave somewhat similarly to a regular two-level system formed by states $\left|e\right\rangle$ and $\left|g\right\rangle$ (see blue and green lines in figure 2(a)). In the second case, the decay of the upper level $\left|f\right\rangle$ is almost instantly followed by the decay of the middle level $\left|e\right\rangle$, so, despite the cascaded process of emission, the two photons are produced nearly simultaneously (the blue line in figure 2(d) is very close to zero, and the green line is close to unity).
Typical transmon circuits have a coupling ratio between the first-excited and second-excited states of $g_{2}(k)/g_{1}(k)\sim\sqrt{3/2}$, see expressions (3.3-3.4) in Koch07 . The corresponding ratio of decay rates is thus $\Gamma_{2}/\Gamma_{1}\sim 1.5$, although some works mention a very similar value of $\Gamma_{2}/\Gamma_{1}\sim 2$ Gasparinetti17 . We plot the probabilities for the common relation of $\Gamma_{2}/\Gamma_{1}=1.5$ in figure 2(b). Albeit, recent progress in superconducting quantum circuits has made possible the creation of the so-called giant-atom transmon with a highly tunable ratio $\beta=\Gamma_{2}/\Gamma_{1}\approx 0.3-60$ Vadiraj21 . However, further down we stick to more common transmon types with ratios of $\Gamma_{2}/\Gamma_{1}=1.5$ and $\Gamma_{2}/\Gamma_{1}=3$.
@@ IV.3 Two-photon spectrum
By using the Fourier transform (7), we can translate the real-space photon amplitudes (10) and (15) into k-space of frequency modes:
@@
where in $\beta(k,t)$ the photon frequency $\omega$ corresponds to the first emitted photon.
Note that if we take $t\rightarrow\infty$, the single-photon amplitude (21), as well as $\alpha(t)$, will be equal to zero. However, the two-photon amplitude (22) for large times takes a non-zero value. This is a logical prediction because the final state of an excited three-level system is obviously two photons in the waveguide with the transmon being in the ground state.
Let’s define the spectral density of spontaneous emission of an initially excited transmon as follows:
@@
We plot this two-photon spectral density in figure 3.
Unlike the probabilities (20), the spectral density (23) depends not only on the decay rates $\Gamma_{1},\Gamma_{2}$, but also on the resonant frequencies of the 3LE. It is typical to describe the energy-level spacing of a transmon via an anharmonicity parameter Koch07 :
@@
where $E_{J}$ is the Josephson energy, and $E_{C}$ is the charging energy — common characteristics of superconducting quantum circuits Blais21 . The subscript $r$ stands for ’relative’, thus expression (24) is dimensionless and shows the relation of energies between states $\left|f\right\rangle,\;\left|e\right\rangle$ and the energies between levels $\left|e\right\rangle,\;\left|g\right\rangle$. In literature it is also common to use a non-relative anharmonicity $\alpha=(\Delta_{\Omega}-\Omega_{1})=\alpha_{r}\Omega_{1}$, which is measured in hertz, but here we choose to use only the relative parameter to avoid confusion with the probability amplitude $\alpha(t)$. Typical experimental values of $\alpha$ usually fall within the range 100-300 MHz Krantz19 , or in relative units from $2\%$ up to $6\%$. Note that the anharmonicity for superconducting qubits is always negative, due to the smaller spacing between the upper levels.
@@
Figure 3: Two-photon spectral density of spontaneous emission from a three-level transmon $S_{sp}(\omega_{1},\omega_{2})\Gamma_{2}^{2}$, plotted using expression (23). (a) is plotted for weak-coupling regime $\Gamma_{2}/\Omega_{1}=0.001$ and typical anharmonicity $\alpha_{r}=-3\%$; Other plots describe stronger coupling for the constant decay rate $\Gamma_{2}/\Omega_{1}=0.02$. (b) and (c) plotted for typical decay rates relation $\Gamma_{2}/\Gamma_{1}=1.5$; (d) and (e) for faster decay of the upper level $\Gamma_{2}/\Gamma_{1}=3$. Plots (b)-(c) share the same colorbar for a more direct comparison of spectral density values, the same applies to plots (d)-(e).
@@
From figure 3(a)-(b) we see that the spectral density of spontaneous emission has two distinct peaks, which are a result of the non-distinguishability of photons. Each peak corresponds to one resonant photon with frequency $\omega=\Omega_{1}$ and a second photon with a smaller frequency $\omega=(1+\alpha_{r})\Omega_{1}$ (remember that $\alpha_{r}$ is negative). But because both photons are indistinguishable, we can’t exactly say which photon is ’first’ and which is ’second’; thus two peaks appear. For bigger anharmonicity, the two peaks are further apart because of the change of transition frequencies in the 3LE. If we implement weak coupling, $\Gamma/\Omega_{1}=0.001$ (in practical circuits it is usually $\Gamma<5$ MHz), both peaks tighten into almost point-like shapes (see figure 3(a)). That means that we have two photons with distinct frequencies, corresponding to atom-transitions, $\Omega_{1}$ and $(1+\alpha_{r})\Omega_{1}$. This behavior is in good agreement with the experimental work by Gasparinetti et al. Gasparinetti17 .
@@
However, if we take stronger coupling, typical for circuit QED systems (for example, $\Gamma_{2}\sim 100$ MHz), we see the appearance of frequency correlation between the two photons. This is expressed by the appearance of a third peak at the resonant frequency $\omega/\Omega_{1}=1$, which starts to become visible for $\Gamma_{2}/\Gamma_{1}=1.5$ in figure 3(c), and is more pronounced for $\Gamma_{2}/\Gamma_{1}=3$ in figure 3(d)-(e). Thus, even if the three-level system has a non-symmetrical level structure, there is a probability that during spontaneous decay the emitted photons will have the same frequency $\omega_{1}=\omega_{2}=\Omega_{1}$. This effect is connected to interference between the first emitted photon and the 3LE being still excited in state $\left|e\right\rangle$, which arises due to the almost equidistant level structure of a low-anharmonicity transmon and rather strong coupling between the artificial atom and the field modes.
The observation of a frequency-correlated photon pair naturally raises the question of whether the two emitted photons are entangled. Entanglement of two photons requires some kind of distinction between different photon states, for example in two polarizations, that can be encoded as two different photon modes. In our work we have only a right moving photon mode, described by photon creation/annihilation operators $a^{\dagger}(x)/a(x)$. Thus, in our derivation we cannot investigate the entanglement of the emitted photon pair properly. However, by considering a bidirectional waveguide with forward-propagating mode $a^{\dagger}(x)$ and backward-propagating mode $b^{\dagger}(x)$ one could investigate the spatial entanglement of two photons, for example in the so-called N00N state $\left(\left|2_{R},0_{L}\right\rangle-\left|0_{R},2_{L}\right\rangle\right)/\sqrt{2}$. We believe that this kind of problem is interesting in itself and warrants its own dedicated research.
@@ IV.4 Emission spectra for identical photons
Figure 4: Spectral density of two identical photons (25) vs. frequency detuning — (a) and (b), and vs. detuning and anharmonicity — (c) and (d). (a) and (c) are plotted for the $\Gamma_{2}/\Gamma_{1}=1.5$ ratio, and (b) and (d) for $\Gamma_{2}/\Gamma_{1}=3$. In (a) and (b), the solid magenta line corresponds to $\alpha_{r}=-3\%$, the green dashed line to $\alpha_{r}=-4\%$, and the black dotted line to $\alpha_{r}=-6\%$. Vertical grey lines show the points for $\delta_{\omega}/\Omega_{1}=\alpha_{r}/2$. In (c) and (d), thin white lines indicate the resonant frequency $\omega=\Omega_{1}$. The decay of the upper level is taken constant $\Gamma_{2}/\Omega_{1}=0.02$.
Let’s study the spectral density (23) for identical photons in closer detail. Taking $\omega_{1}=\omega_{2}$ we get:
@@
where we introduce the detuning from the resonant frequency for both photons as $\delta_{\omega}=\omega-\Omega_{1}$. We can interpret function (25) as a diagonal cross-section from two-photon plots from figure 3.
The function (25) has two maxima: the first one corresponds to the frequency between the lower levels $\Omega_{1}$ and thus $\delta_{\omega}=0$, and the second one corresponds to the decay of the upper level with an anharmonicity shift, $\delta_{\omega}=-\alpha_{r}\Omega_{1}/2$. Both these peaks can be observed in figure 4(a)-(b), for the decay rates $\Gamma_{2}/\Gamma_{1}=1.5$ and $\Gamma_{2}/\Gamma_{1}=3$. A two-dimensional plot of (25), showing how the spectral density also changes with anharmonicity, is presented in figure 4(c)-(d).
The maximum of the spectral density for identical photons should be for $\alpha_{r}\rightarrow 0$, i.e., in the case of equidistant levels with $\Omega_{2}=2\Omega_{1}$. However, as can be seen from figure 4, the frequencies of the two photons do not always have to be resonant with $\Omega_{1}$. For example, if the ratio between the decay rates is $\Gamma_{2}/\Gamma_{1}=1.5$, like in figure 4(a), the emitted photons both would have the maximum spectral density at a frequency close to $\Omega_{1}(1+\alpha_{r}/2)$. Thus, by changing both the decay rates $\Gamma_{1},\Gamma_{2}$ and the anharmonicity parameter $\alpha_{r}$ of the excited three-level transmon, it is possible to create correlated photon pairs with the same frequencies. Superconducting circuits are very well suited for this task, since they allow tuning of the couplings and anharmonicity parameters either by sample production techniques or via the application of external magnetic fields.